A 54.5 kg skier is moving at 11.4 m/s on a frictionless horizontal, snow covered plateau when she encounters a rough patch 1.1 m long. The coefficient of kinetic friction between this patch and her skis is 0.24. After crossing the rough patch, what is the speed of the skier in ms-1 ? g=9.8ms-2
1
Expert's answer
2018-11-05T15:47:09-0500

change of energy=work done

(mv_f^2)/2-(mv_i^2)/2=-F_frict l

(mv_f^2)/2-(mv_i^2)/2=-μmgl

So

v_f=√(v_i^2-2μgl)=√(〖11.4〗^2-2×0.24×9.8×1.1)=11.2 m/s